OpenStudy (anonymous):

How to prove or explain why sin(2x) does not equal 2sin(x) ?

OpenStudy (aravindg):

sin 2x =sin (x+x) use sin(A+B)

OpenStudy (anonymous):

how would I use the sin(A+B)=sinacosb+cosasinb formula to solve that?

OpenStudy (aravindg):

here x=A=B

OpenStudy (aravindg):

just use that !

OpenStudy (aravindg):

sin(A+B)=sinacosb+cosasinb A=B=x

OpenStudy (anonymous):

So what would A and B be in the equation? 2?

OpenStudy (anonymous):

"How to prove or explain why sin(2x) does not equal 2sin(x) ? " Find a counter example.

OpenStudy (aravindg):

A=B=x

OpenStudy (anonymous):

sin(2x) is a horizontal stretch of factor 0.5 2sin(x) is a vertical stretch of factor 2

OpenStudy (e.mccormick):

Another way to prove negatives is with an example case that fails. \(2\sin{\frac{\pi}{2}}\ne\sin\pi\)

OpenStudy (anonymous):

One good counter example is \(x=90^\circ\).

OpenStudy (aravindg):

sin x cos x+cos x sin x=2 sinx cos x is not equal to 2 sin x